OpenStudy (anonymous):

Find the critical points of f(x)=(2^x)(sin(x) . I can only find -.96, but my online homework is telling me there are two more. Please explain

OpenStudy (amistre64):

f' = 2^x cos(x) + 2^x ln(2) sin(x) = 2^x ( cos(x) + ln(2) sin(x)) that doesnt appear to be undefined so it should amount to the zeros right?

OpenStudy (jamesj):

Yes, in fact there are infinitely many solutions, as f'(x) = 0 when tan x = -1/ln 2

OpenStudy (amistre64):

yeah, those infinite ones have some extra jargon in them ... ;)

OpenStudy (jamesj):

And we should expect an infinite number of critical points, as this will be an oscillating function where the amplitude increases in the positive x direction.

OpenStudy (jamesj):

In physics, functions like this come up all the time.

OpenStudy (anonymous):

how do i find the solutions within the interval [-2,6]?

OpenStudy (amistre64):

by finding out when the derivative f'(x)= 0 between -2 and 6, and check the end points as well

OpenStudy (amistre64):

since 2^x never equals zero, we only have to consider the rest of it: cos(x) + ln(2) sin(x) = 0 ln(2) sin(x) = -cos(x) -ln(2) tan(x) = 1 and 1 = -cot(x)/ln(2) maybe....

OpenStudy (anonymous):

how do i solve those for x? inverse cot(x)?